What are the equilibrium values of C,I and r?

GDP (Y) i=5,000

C=740+0.5 (Y-T)


I=1500 - 50r where r is the real rate of interest in percent

T= 1259

G=1125


1
Expert's answer
2022-02-13T11:48:25-0500

In equilibrium Y = C + I + G, so:

"740 + 0.5\u00d7(5,000 - 1,259) + 1500 - 50r + 1,125 = 5,000,"

50r = 235.5,

r = 4.71%,

C = 2,610.5,

I = 1500 - 50×4.71 = 1,264.5.